OpenStudy (anonymous):

Which triangle side lengths form a right triangle? Choose all that apply. 5, 6, 7 6, 8, 14 8, 15, 17 10, 12, 15 15, 20, 25 16, 18, 20

Nnesha (nnesha):

apply Pythagorean theorem \[\huge\rm a^2 +b^2=c^2\] replace a b c by their values if both sides are equal then yes thse 3 numbers form a right triangle

Nnesha (nnesha):

remember : c is the longest side of right triangle so for first one c=7 plug in\[\huge\rm 5^2 +6^2 =7^2\] both sides are equal ?

OpenStudy (anonymous):

no both sides arent equal right?

OpenStudy (anonymous):

so which one would be the best reponces? @Nnesha

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @eugeniiaxo no both sides arent equal right? \(\color{blue}{\text{End of Quote}}\) yes both sides are not equal so that's wrong

Nnesha (nnesha):

try other numbers

Nnesha (nnesha):

if you get equal sides then that would be your answer

OpenStudy (anonymous):

15, 20, 25 this is one

Nnesha (nnesha):

yes right that would form a right traingle

OpenStudy (anonymous):

is that the one one that works cause thats the only on i got

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

there is one more

OpenStudy (anonymous):

&& 8 15 17 right

Nnesha (nnesha):

yep :=)